# Automatic Database Configuration Debugging using Retrieval-Augmented Language Models

### v1

@@ 6.1. Experimental Setup
Datasets. We use four real-world datasets in two types of experimental settings (NL and Runnable),. Table 1 provides statistics of the datasets. Note that, the training labels can be automatically annotated. For the original training data, we use regular expressions to extract relevant knobs from accepted answers or comments on questions from StackOverflow or the MySQL Forum. For the augmented training data, we generate questions based on a knob’s manual, with the knob serving as the label corresponding to the questions. To validate labeling quality, we have randomly sampled some training and testing cases for manual verification, which have confirmed that our automatic labeling is accurate.
@@ 6.2. Overall Comparisons
Table 2. Overall results, where we measure Precision, Recall and F1-score in NL setting, and SuccessRate in Runnable setting.
@@
Evaluation in the NL Setting. We report the Recall, Precision and F1-scores of different approaches on the three datasets, MySQL SO, MySQL Forum and PG SO, in the NL setting, as shown in Table 2. The PLM-based method does not perform well due to the limited availability of training data. Even though PLM+DA incorporates more training data, the improvement still remains marginal. This is because the PLM, with its fewer parameters, struggles to understand the context, particularly in longer and expert-domain texts.
LLM-based methods (LLMs in Table 2) perform better than PLM-based methods. However, they also have unsatisfactory performance. The main reason is hallucination - they often predict knobs that do not exist in reality, even for GPT-4. Moreover, GPT models outperform open-source LLMs, while smaller-sized models (Llama8B and Qwen7B) fail to deliver satisfactory performance.
This phenomenon gets better with LLMs (all knobs), because this essentially turns a generation problem into a selection problem, which significantly reduces hallucinations. However, these methods still have limitations. In the database configuration debugging scenario, the database contains a large number of knobs with complex functions, and there are intricate logical relationships between the knobs and anomalies. To deal with such complex questions, LLM models still lack expert knowledge and experience. Additionally, we find that sometimes GPT-4 does not perform better than GPT-3.5, such as in the MySQL Forum, because GPT-4 may not strictly output results as our required format.
@@
In detailed analysis, the number of knobs predicted on average by Andromeda is $4.85$ with GPT-3.5 and $2.63$ with GPT-4, which means that users can solve most problems by adjusting just a few knobs. Interestingly, Andromeda achieves higher accuracy when using GPT-3.5 compared to GPT-4. This is partly attributed to GPT-4’s tendency to deviate from our specified output requirements, and it produces fewer knobs In contrast, when the number of adjustable knobs is controlled, GPT-3.5 recommends more knobs, making it easier to find knobs that can solve the problem. Compared with other methods, Andromeda shows significant improvements compared with GPTs and GPTs (all knobs).
@@ 6.3. Evaluation on Document Retrieval
Exp-2: How does data augmentation work in document retrieval module? In order to see the effect of the augmentation module, we consider an Andromeda-docs-only version of our framework by disabling telemetry analysis. We compare Andromeda-docs-only with the following document augmentation methods: (1) No Augment: training model using the original training data (as shown in Table 1). (2) CoT Augment: let LLM generate the same volume questions from manuals using CoT method. The results are provided in Table 3 for Document retrieval recall and Table 4 for F1-score, which show that Andromeda-docs-only outperforms these two methods. Andromeda-docs-only and CoT Augment perform better than No Augment, showing that the new generated question-knob pairs are helpful. Andromeda-docs-only performs better than CoT Augment due to our proposed logic chain strategy.
Finding 2: The logic chain based data synthesis strategy for training data augmentation is very helpful and outperforms other data augmentation methods.
@@
Exp-4: How does inaccuracies in the retrieval sources impact the performance of Andromeda? We manually inject noise into these sources by replacing the ground-truth labels in historical questions and database manuals with incorrect ones in various ratios, i.e., 0.1, 0.2, 0.3, 0.4, and 0.5. Figure 7 shows that as the noise ratio increases, the overall performance of Andromeda decreases, highlighting the importance of the quality of retrieval sources. In particular, on some datasets (e.g., MySQL SO in the Noisy Manuals setting), the performance does not significantly change. This is because, despite the presence of noisy documents, useful ones can still be retrieved to guide the LLMs.
@@ 6.4. Evaluation on Telemetry Data Analysis
The results are presented in Table 7. GPTs and Andromeda-telemetry-only perform similarly. This indicates that using telemetry data analysis alone has limit effects as it only diagnosis a part of configuration issues that are highly related to telemetries: the telemetry data analysis can provide description on anomalous telemetry changes, which are useful for diagnosing anomalies caused by these anomalous changes. However, some questions are not due to the state of runtime environment resources, so using only telemetry data analysis is insufficient.
@@ 6.5. Efficiency Evaluation
Exp-7: Whether Andromeda effectively reduces the end-to-end time? We evaluate whether Andromeda effectively reduce the end-to-end time (i.e., operational burdens) for DBAs To this end, we employ two experienced DBAs to solve questions in the Runnable setting and provide them with historical questions, database manuals, and telemetry data. We then consider three strategies: (1) using manuals: DBAs manually search DBMS manuals to provide diagnostic suggestions for tuning knobs; (2) using search engines: DBAs use search engine to provide diagnostic suggestions for tuning knobs; (3) using Andromeda: DBAs follow Andromeda ’s recommendations to address the issues. We compare the average time taken by the DBAs to troubleshoot configuration issues, as well as the SuccessRate scores. As shown in Table 9, Andromeda achieves a higher SuccessRate than the alternative strategies and significantly reduces end-to-end time. These results demonstrate that Andromeda effectively reduces end-to-end time for DBAs while providing accurate troubleshooting suggestions.
Table 9. Comparison of on SuccessRate and end-to-end time with using manuals and search engines in Runnable setting.

### v2

@@ 6.1. Experimental Setup
Datasets. We use four real-world datasets in two types of experimental settings (NL and Runnable). Table 1 provides statistics of the datasets. Note that, the training labels can be automatically annotated. For the original training data, we use regular expressions to extract relevant knobs from accepted answers or comments on questions from StackOverflow or the MySQL Forum. For the augmented training data, we generate questions based on a knob’s manual, with the knob serving as the label corresponding to the questions. To validate labeling quality, we have randomly sampled some training and testing cases for manual verification, which have confirmed that our automatic labeling is accurate.
@@ 6.2. Overall Comparisons
Table 2. Overall results, where we measure Precision, Recall and F1-score in the NL setting, and SuccessRate in the Runnable setting.
@@
Evaluation in the NL Setting. We report the Recall, Precision and F1-scores of different approaches on the three datasets, MySQL SO, MySQL Forum and PG SO, in the NL setting, as shown in Table 2.
The PLM-based method does not perform well due to the limited availability of training data. Even though PLM+DA incorporates more training data, the improvement still remains marginal. This is because the PLM, with its fewer parameters, struggles to understand the context, particularly in longer and expert-domain texts.
LLM-based methods (LLMs in Table 2) perform better than PLM-based methods. However, they also have unsatisfactory performance. The main reason is hallucination: they often predict knobs that do not exist in reality, even for GPT-4. Moreover, GPT models outperform open-source LLMs, while smaller-sized models (Llama8B and Qwen7B) fail to deliver satisfactory performance.
This phenomenon gets better with LLMs (all knobs), because this essentially turns a generation problem into a selection problem, which significantly reduces hallucinations. However, these methods still have limitations. In the database configuration debugging scenario, the database contains a large number of knobs with complex functions, and there are intricate logical relationships between the knobs and anomalies. To deal with such complex questions, LLM models still lack expert knowledge and experience. Additionally, we find that sometimes GPT-4 does not perform better than GPT-3.5 (e.g., in the MySQL Forum), because GPT-4 may not strictly output results as our required format.
@@
In detailed analysis, the number of knobs predicted on average by Andromeda is $4.85$ with GPT-3.5 and $2.63$ with GPT-4, which means that users can solve most problems by adjusting just a few knobs. Interestingly, Andromeda achieves higher accuracy when using GPT-3.5 compared to GPT-4. This is partly attributed to GPT-4’s tendency to deviate from our specified output requirements, and it produces fewer knobs. In contrast, when the number of adjustable knobs is controlled, GPT-3.5 recommends more knobs, making it easier to find knobs that can solve the problem. Compared with other methods, Andromeda shows significant improvements compared with GPTs and GPTs (all knobs).
@@ 6.3. Evaluation on Document Retrieval
Exp-2: How does data augmentation work in document retrieval module? In order to see the effect of the augmentation module, we consider an Andromeda-docs-only version of our framework by disabling telemetry analysis. We compare Andromeda-docs-only with the following document augmentation methods: (1) No Augment: training model using the original training data (as shown in Table 1). (2) CoT Augment: let LLM generate the same volume questions from manuals using the CoT method. The results are provided in Table 3 for Document retrieval recall and Table 4 for F1-score, which show that Andromeda-docs-only outperforms these two methods. Andromeda-docs-only and CoT Augment perform better than No Augment, showing that the new generated question-knob pairs are helpful. Andromeda-docs-only performs better than CoT Augment due to our proposed logic-chain strategy.
Finding 2: The logic-chain based data synthesis strategy for training data augmentation is very helpful and outperforms other data augmentation methods.
@@
Exp-4: How do inaccuracies in the retrieval sources impact the performance of Andromeda? We manually inject noise into these sources by replacing the ground-truth labels in historical questions and database manuals with incorrect ones in various ratios, i.e., 0.1, 0.2, 0.3, 0.4, and 0.5. Figure 7 shows that as the noise ratio increases, the overall performance of Andromeda decreases, highlighting the importance of the quality of retrieval sources. In particular, on some datasets (e.g., MySQL SO in the Noisy Manuals setting), the performance does not significantly change. This is because, despite the presence of noisy documents, useful ones can still be retrieved to guide the LLMs.
@@ 6.4. Evaluation on Telemetry Data Analysis
The results are presented in Table 7. GPTs and Andromeda-telemetry-only perform similarly. This indicates that using telemetry data analysis alone has limit effects as it only identifies a part of configuration issues that are highly related to telemetries: the telemetry data analysis can provide description on anomalous telemetry changes, which are useful for diagnosing anomalies caused by these anomalous changes. However, some questions are not due to the state of runtime environment resources, and thus using only telemetry data analysis is insufficient.
@@ 6.5. Efficiency Evaluation
Exp-7: Whether Andromeda effectively reduces the end-to-end time? We evaluate whether Andromeda effectively reduces the end-to-end time (i.e., operational burdens) for DBAs. To this end, we employ two experienced DBAs to solve questions in the Runnable setting and provide them with historical questions, database manuals, and telemetry data. We then consider three strategies: (1) using manuals: DBAs manually search DBMS manuals to provide diagnostic suggestions for tuning knobs; (2) using search engines: DBAs use search engine to provide diagnostic suggestions for tuning knobs; (3) using Andromeda: DBAs follow Andromeda’s recommendations to address the issues. We compare the average time taken by the DBAs to troubleshoot configuration issues, as well as the SuccessRate scores. As shown in Table 9, Andromeda achieves a higher SuccessRate than the alternative strategies and significantly reduces end-to-end time. These results demonstrate that Andromeda effectively reduces end-to-end time for DBAs while providing accurate troubleshooting suggestions.
Table 9. Comparison on SuccessRate and end-to-end time of using Andromeda with using manuals and using search engines in Runnable setting.
@@ 7. Conclusions and Future Work
Acknowledgements.
We thank the anonymous reviewers for their constructive feedback. This work was partly supported by the NSF of China (62436010 and 62441230), Alibaba Innovative Research Program, the Beijing Natural Science Foundation (L222006), the Research Funds of Renmin University of China, the Outstanding Innovative Talents Cultivation Funded Programs 2023 of Renmin University of China and Guangzhou Municipality Big Data Intelligence Key Lab (2023A03J0012).